OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = −three fourthsx + 8 y = −three fourthsx + 3

OpenStudy (anonymous):

Hei, friend, it hould be you give the question in Mathematics subject, but it is no problem. I will answer it :) First, find the gradien 4x − 3y = −12 y = (-4x -12)/3 y = -4x/3 4 It means that the gradien is -4/3 because it is parallel use the formula to find the equation y-y1 = m(x - x1) y - 4 = -4/3 (x +3) y = -4x/3 So, the equation is y = -4x/3